If a,b,c areunit vectors and a+b+c=i+j-2k,then a.b+b.c+c.a=

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

a+b+c=i+j-2k

|a+b+c|2=|i+j-2k|2

|a| +|b| +|c|    +2(a.b +b.c+c.a)  =6

 2(a.b +b.c+c.a)  =6 -3 =3

(a.b +b.c+c.a)  =3/2

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin